OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. A. y = 3x - 10 B. y = 3x - 6 C. y = 3x - 2 D. y = 3x - 12

OpenStudy (anonymous):

since the line is parallel to given line..their slopes are equal..can you find the slope from the given line??

OpenStudy (anonymous):

@abhi7 is it d?

OpenStudy (anonymous):

are u just guessing or u have calculated??

OpenStudy (anonymous):

honestly just guesssing, i don't know where to even begin on this.

OpenStudy (anonymous):

compare the given equation y=3x-6 to y=mx+c..what is m here??

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yes here m is the slope of the line y=3x-6..since the line we need to find is parallel to this line its slope will also be 3

OpenStudy (anonymous):

now using point slope form y=mx+c..since (2,-4) passes through the line and m=3..use (x,y)=(2,-4) and m=3 in y=mx+c to find c..try to find c now

OpenStudy (anonymous):

y= 3 2 + -4 y= 6 + -4 y=2 so its y = 3x - 2?

OpenStudy (anonymous):

you have put -4 in c..-4 should be put in y

OpenStudy (anonymous):

find the value of c.. we have given (x,y)..try one more time

OpenStudy (anonymous):

- 4 = 3.2 + c? is that the right equation?

OpenStudy (anonymous):

yes now find c

OpenStudy (anonymous):

-4 = 6 -6 -6 c= -10? oh my god im sorry... thanks for the patience.

OpenStudy (anonymous):

correct..after finding c..you have slope(m) and y-intercept(c) of the given line..substitute these values in y=mx+c and you have the required equation..

OpenStudy (anonymous):

Thank you so much!(: